# PrivATE: Differentially Private Confidence Intervals for Average Treatment Effects

### v1

@@ 1 Differentially private ATE estimation
In the first step, we privatize the original ATE estimate $\hat{\tau}$. Directly employing Definition 3.2 requires calculating the model sensitivity, which can be computationally expensive or even infeasible. We instead exploit the fact that a suitably scaled gross-error sensitivity $\gamma$ can upper bound the global sensitivity [5]. Therefore, we first derive the influence function of the AIPW estimator so that we are then able to privatize $\hat{\tau}$ based on $\gamma$.
For the AIPW estimator, the influence function score is given by [54]
$\Gamma_{\hat{\eta}}(z)=\left(\frac{a_{i}}{\hat{\pi}(x_{i})}-\frac{1-a_{i}}{1-\hat{\pi}(x_{i})}\right)y_{i}-\frac{\left(1-\hat{\pi}(x_{i})\right)\hat{\mu}_{1}(x_{i})+\hat{\pi}(x_{i})\hat{\mu}_{0}(x_{i})}{\hat{\pi}(x)\left(1-\hat{\pi}(x_{i})\right)}(a_{i}-\hat{\pi}(x_{i})).$
Since the AIPW estimator can equivalently be written as $\hat{\tau}=\sum_{i=1}^{n}\Gamma_{\hat{\eta}}(z_{i}),$ the IF of the AIPW estimator directly follows via [^5] For an introduction to the construction of IFs for common statistics such as the mean, we refer to [33].
@@
Let $z:=(a,x,y)$ define a data sample following the joint distribution $\mathcal{Z}$, and let $\hat{\eta}=(\hat{\mu},\hat{\pi})$ be the estimated nuisance functions. Furthermore, let $D$ be the training dataset with $|D|$ = $n$. For $U\sim\mathcal{N}(0,1)$, $(\varepsilon,\delta)$-DP is fulfilled by
@@ 2 Differentially private variance estimation
The variance estimate depends on the estimated nuisance functions. Standard variance estimators are, in general, only consistent if both nuisance functions are correctly specified [28]. Thus, straightforward estimators, such as the influence function-based estimator  [18], are not doubly robust and, therefore, can easily lead to undercoverage of the CIs in practice, which is undesirable in safety-critical applications. As a remedy, we estimate $\hat{\sigma}_{\textrm{{AIPW}}}^{2}$ by a doubly robust variance estimator.
It is common to obtain such an estimator through nonparametric bootstrapping [27, 62]. However, bootstrapping requires many queries to the data, which is disadvantageous for estimation under DP guarantees, as each query requires a separate privacy budget.
Instead, in our work, we estimate $\hat{\sigma}_{\textrm{{AIPW}}}^{2}$ via the empirical sandwich estimator [36]. This has the following benefits: (i) the empirical sandwich estimator is doubly robust, and (ii) it consistently estimates the variance of the AIPW estimator [62]. Importantly, for the AIPW estimator, the sandwich estimator can be directly represented by the IF (see Supplement A)
@@
Additionally, we observe that the differentially private variance is naturally positive. To impose this restriction during the privatization process, we employ a truncated privacy mechanism: Let $\mathbf{f}:\mathcal{X}^{d}\mapsto\mathbb{R}^{d}$ and let $\Delta_{2}(\mathbf{f})=\sup_{D\sim D^{{}^{\prime}}\!\!,\,\mathbf{x}\in\mathcal{X}^{d}}||\mathbf{f}_{D}(\mathbf{x})-\mathbf{f}_{D^{{}^{\prime}}}(\mathbf{x})||$ denote the $l_{2}$-sensitivity of $\mathbf{f}$. Furthermore, let $\mathbf{U}\sim\mathcal{N}(0,\sigma^{2})$ for $\sigma\geq\frac{1}{\varepsilon}\sqrt{2\ln{(1.25/\delta)}}\cdot\Delta_{2}(\mathbf{f})$. Then,
$\mathbf{f}^{\textrm{{DP}}}_{D}(\mathbf{x})=\max\{0,\mathbf{f}_{D}(\mathbf{x})+\mathbf{U}\}$ (10)
@@
Let $z:=(a,x,y)$ define a data sample following the joint distribution $\mathcal{Z}$, and let $\hat{\eta}=(\hat{\mu},\hat{\pi})$ be the estimated nuisance functions. Furthermore, let $D$ be the training dataset with $|D|$ = $n$. For $U\sim\mathcal{N}(0,1)$, we define
@@ 3 Constructing CIs for the ATE under DP
One might think that it is sufficient to simply privatize the variance to ensure DP. However, that is not the case: In the following, we clarify that achieving DP for the variance alone would indeed be sufficient—but only in the asymptotic regime, due to asymptotic normality. However, this leads to problematic implications in finite samples. Specifically, the privatized variance may underestimate the true variance, leading to undercoverage in the CIs.
Observe that the AIPW estimator $\hat{\tau}$ is a consistent estimator for the true ATE $\tau_{0}$ and asymptotically normally distributed with variance $\sigma^{2}_{\textrm{{AIPW}}}$ [e.g., 64], i.e.,
@@
In finite settings, however, the privatization introduces additional noise on $\hat{\tau}$ (and thus variance), which we account for in the following when constructing our CIs. This is in line with prior literature postulating that DP inference should be conservative to ensure reliability even in small sample sizes [e.g., 39]. Observe that $(\hat{\tau}_{\textrm{{DP}}}-\hat{\tau})=:B$ is itself a random variable (for fixed $\hat{\tau}$) following distribution $\mathcal{N}\left(0,r(\varepsilon,\delta,n)^{2}\right)$, where
@@
To incorporate the finite sample variance stemming from the privatization, we see from (14) that, for the variance $V$, it must hold that: $V=\sigma_{\textrm{{AIPW}}}^{2}+nr(\varepsilon,\delta,n)^{2}$. ropose a truncated Gaussian noise privacy mechanism. For ease of notation, we define the truncated Gaussian distribution as $\mathcal{N}^{+}(\mu,\sigma)$ with density
@@
Let $z:=(a,x,y)$ define a data sample following the joint distribution $\mathcal{Z}$, and let $\hat{\eta}=(\hat{\mu},\hat{\pi})$ be the estimated nuisance functions. We define the privacy budget $(\varepsilon,\delta)$ and the sub-budgets $\varepsilon_{1},\varepsilon_{2},\delta_{1},\delta_{2}$, for ATE and variance estimation under DP such that $\varepsilon_{1}+\varepsilon_{2}=\varepsilon$ and $\delta_{1}+\delta_{2}=\delta$. Furthermore, let $D$ be the training dataset with $|D|$ = $n$, and let $\hat{\tau}_{\textrm{{DP}}}$ denote the $(\varepsilon_{1},\delta_{1})$-differentially private ATE estimate and $\hat{\sigma}_{\textrm{{DP}}}$ denote the $(\varepsilon_{2},\delta_{2})$-differentially private variance estimate. Then, the valid and $(\varepsilon,\delta)$-differentially private CI is given by
@@ 5 Experiments
Performance metrics: We evaluate PrivATE based on the notion of faithfulness [e.g., 59], i.e., we compute the empirical coverage of the CIs and assess whether it surpasses the $1-\alpha$ threshold.
@@ 5.1 Evaluation on synthetic datasets
Performance across varying privacy budgets: We now vary the privacy budget $\varepsilon$ in Figure 4. We observe that PrivATE behaves as expected: for increasing privacy budgets, the intervals become tighter, eventually converging to the non-private estimate. We note that privacy inevitably induces a trade-off regarding the width of the interval, which we discuss in Supplement D.
Performance across varying sample sizes: We further analyze the performance of PrivATE across different sample sizes $n$ (see Fig. 4). The results are as expected and confirm our theoretical insights: For larger sample sizes, the CIs become tighter, converging to the asymptotic solution. Additionally, we observe that PrivATE is not sensitive to the choice of the base learner: both versions with (i) kernel ridge regression (=Kernel) and (ii) neural network (=NN) lead to similar results. This again demonstrates the flexibility of our framework.
@@ 5.2 Evaluation on medical datasets
We demonstrate the real-world applicability of PrivATE in a medical context on the Right Heart Catheterization (RHC) dataset [16]. The dataset contains observations of 5735 patients admitted to intensive care units. We aim to predict the ATE of receiving a Swan-Ganz catheterization on the probability of death following hospitalization. Further details on the RHC dataset are in Supplement C.
@@
Results: In Table 2, we present the CIs returned by PrivATE compared to an adapted version of the naïve CIs constructed with a valid variance, which includes the correction term for the finite-sample variance stemming from ATE privatization. Again, we do not aim to benchmark the methods as the naïve CIs are non-private. Rather, we aim to understand whether the CIs from PrivATE are faithful. We observe that the PrivATE CIs include the valid, naïve CIs, which we see as an indication of validity. In other words, our PrivATE gives CIs that, in addition to fulfilling $(\varepsilon,\delta)$-DP, remain informative.
Medical interpretation: We observe that we always receive a positive private treatment effect $\hat{\tau}_{\textrm{{DP}}}$, indicating an average increase in the probability of death after receiving a Swan-Ganz catheter. The negative effect on survival is likely to be attributed to the life-threatening complications and risks of introducing the catheter, such as pulmonary artery rupture [e.g., 41]. The positive $\hat{\tau}_{\textrm{{DP}}}$ also coincides with the non-private ATE estimate $\hat{\tau}=1.17$ and is in line with findings in the literature [e.g., 16, 58].
Extension to interventional RCT data: Our PrivATE framework applies to both observational and experimental data. In experimental randomized controlled trials (RCTs), the propensity score is known and thus does not need to be estimated. Note that in some cases, however, estimating the propensity score, even in RCTs, can be beneficial [63]. So far, we have evaluated PrivATE on observational data. To show the applicability of PrivATE to RCTs, we conduct additional experiments in Supplement E.
@@ A.1 Estimating the average treatment effect
Under the three standard assumptions of causal inference, i.e., positivity, consistency, and unconfoundedness [e.g., 17, 57], the ATE is point-identified as
@@
However, estimating $\hat{\tau}$ via the empirical average of the stated difference is only unbiased if the treatment assignment is randomized and if it does not depend on $X$ (i.e., if for a constant $\pi(x)$). To account for a non-random treatment assignment in observational data, the inverse propensity weighted (IPW) estimator constructs $\hat{\tau}$ by estimating $\hat{\pi}$ in the first step and then employing the estimates to take the propensity weighted empirical average of the outcomes [56].
Yet, misspecification of $\hat{\pi}$ or $\hat{\mu}$ can lead to biased ATE estimates. As a remedy, the augmented inverse propensity weighted (AIPW) estimator [55] employs an orthogonal loss to fulfill the double robustness property. Doubly robust estimators are consistent if either $\hat{\mu}$ or $\hat{\pi}$ are correctly specified. The AIPW estimator is then given by
@@
Hence, in contrast to the IPW estimator, the AIPW estimator is an unbiased estimator of ATE and is asymptotically normal distributed [64]. This is important as we (i) aim to consistently estimate the ATE, and (ii) can derive valid CIs due to the asymptotic normality to the estimator. In our work, we thus aim to derive differentially private CIs for the ATE estimated through the AIPW estimator.
@@ A.2 Differential privacy
In our work, we employ output perturbation based on the Gaussian mechanism. Of note, output perturbation is especially suitable for our task of constructing CIs for the estimated ATE due to our three reasons: output perturbation (i) retains the adherence to the causal assumptions, (ii) is model agnostic, and (iii) preserves the ability of the AIPW estimator to address the fundamental problem of causal inference [60]. In contrast, other approaches (i.e., input, objective, and gradient perturbations) fail to fulfill at least one of the points (i)–(iii) and would thus be not suitable for estimating the ATE with DP guarantees in a model-agnostic way.
@@
Lemma A.1 (Sequential composition property [23]).
@@
Lemma A.2 (Post-processing property [20]).
@@ A.3 Extended related work
Uncertainty quantification for causal quantities: Many different approaches for quantifying uncertainty in causal estimates exist in the literature. Many works rely on Bayesian methods [e.g., 3, 32, 38, 35]. However, Bayesian methods require prior distributions informed by domain knowledge. Therefore, they are less robust to model misspecification and, therefore, not suitable for our model-agnostic framework. A different stream focuses on estimating the conditional distribution of the treatment effect, which then can be employed for bootstrapping to construct uncertainty intervals [46, 47, 48]. Another stream of literature provides finite-sample uncertainty guarantees based on conformal prediction methods [e.g., 45, 59]. These methods can only construct intervals for the potential outcomes, not the treatment effects. Although one could construct from conformal prediction intervals for treatment effects based on the potential outcomes, those intervals commonly tend to be very wide and, therefore, less informative in practice. Importantly, none of the works above can provide $(\varepsilon,\delta)$-differentially private confidence intervals, which is the novelty of our work.
@@ A.4 Doubly robust variance estimation
Standard variance estimators for the AIPW estimator are, in general, only consistent if both nuisance functions are correctly specified [28]. Thus, these estimators are not doubly robust and, therefore, can easily lead to undercoverage of the CIs in practice, a highly undesirable scenario in safety-critical applications. We thus aim to estimate $\hat{\sigma}_{\textrm{{AIPW}}}^{2}$ through a doubly robust variance estimator.
A common such estimator is built upon nonparametric bootstrapping [27, 62]. However, bootstrapping requires many queries to the training data, which is highly disadvantageous for estimating the variance with DP guarantees. Further, bootstrapping is computationally expensive. This is unlike doubly robust variance estimators, which have been proposed only for specific applications, namely, survey data with missing data imputation [e.g., 30, 42].
In our work, we estimate $\hat{\sigma}_{\textrm{{AIPW}}}^{2}$ via the empirical sandwich estimator [36]. This has the following benefits: (i) the empirical sandwich estimator is doubly robust, and (ii) it consistently estimates the variance of the AIPW estimator [62].
@@
Define $\mathbf{B}(\theta):=\mathbb{E}[-\frac{\partial}{\partial\theta}\psi(Z_{i},\theta)]$ with $\mathbf{B}_{n}(\hat{\theta}):=\frac{1}{n}\sum_{i=1}^{n}[-\frac{\partial}{\partial\hat{\theta}}\psi(Z_{i},\hat{\theta})]$ and $\mathbf{M}(\theta):=\mathbb{E}[\psi(Z_{i},\theta)\psi(Z_{i},\theta)^{T}]$ with $\mathbf{M}_{n}(\hat{\theta}):=\frac{1}{n}\sum_{i=1}^{n}[\psi(Z_{i},\hat{\theta})\psi(Z_{i},\hat{\theta})^{T}]$. Then, the variance of $\hat{\theta}$ can be estimated by the empirical sandwich variance estimator $\mathbf{V}_{n}(\hat{\theta})$ as

### v2

@@ 1 Differentially private ATE estimation
In the first step, we privatize the original ATE estimate $\hat{\tau}$. Directly employing Definition 3.2 requires calculating the model sensitivity, which can be computationally expensive or even infeasible. We instead exploit the fact that a suitably scaled gross-error sensitivity $\gamma$ can upper bound the global sensitivity (Avella-Medina, 2021). Therefore, we first derive the IF of the AIPW estimator so that we are then able to privatize $\hat{\tau}$ based on $\gamma$.
For the AIPW estimator, the influence function score is given by (Robins and Rotnitzky, 1995; Wager, 2024)
$\Gamma_{\hat{\eta}}(z)=$ $\left(\frac{a_{i}}{\hat{\pi}(x_{i})}-\frac{1-a_{i}}{1-\hat{\pi}(x_{i})}\right)y_{i}$ (7)
$-\frac{\left(1-\hat{\pi}(x_{i})\right)\hat{\mu}_{1}(x_{i})+\hat{\pi}(x_{i})\hat{\mu}_{0}(x_{i})}{\hat{\pi}(x)\left(1-\hat{\pi}(x_{i})\right)}(a_{i}-\hat{\pi}(x_{i})).$
Since the AIPW estimator can equivalently be written as $\hat{\tau}=\sum_{i=1}^{n}\Gamma_{\hat{\eta}}(z_{i})$, the IF of the AIPW estimator directly follows via [^4] For an introduction to the construction of IFs for common statistics such as the mean, we refer to (Hines et al., 2022).
@@
[…]
@@ 2 Differentially private variance estimation
The variance estimate depends on the estimated nuisance functions. Standard variance estimators are, in general, only consistent if both nuisance functions are correctly specified (Gruber and van der Laan, 2012). Thus, straightforward estimators, such as the influence function-based estimator  (Deville, 1999), are not doubly robust and, therefore, can easily lead to undercoverage of the CIs in practice, which is undesirable in safety-critical applications. As a remedy, we estimate $\hat{\sigma}_{\textrm{{AIPW}}}^{2}$ by a doubly robust variance estimator.
It is common to obtain such an estimator through nonparametric bootstrapping (Funk et al., 2011; Shook-Sa et al., 2024). However, bootstrapping requires many queries to the data, which is disadvantageous for estimation under DP guarantees, as each query requires a separate privacy budget.
Instead, in our work, we estimate $\hat{\sigma}_{\textrm{{AIPW}}}^{2}$ via the empirical sandwich estimator (Huber, 1967). This has the following benefits: (i) the empirical sandwich estimator is doubly robust, and (ii) it consistently estimates the variance of the AIPW estimator (Shook-Sa et al., 2024). Importantly, for the AIPW estimator, the sandwich estimator can be directly represented by the IF (see Supplement A)
@@
[…]
$\mathbf{f}^{\textrm{{DP}}}_{D}=\max\{0,\mathbf{f}_{D}+\mathbf{U}\}$ (12)
@@
[…]
$\hat{\sigma}^{2}_{\textrm{{DP}}}:=$ $\max\bigg\{0,\hat{\sigma}_{\textrm{{AIPW}}}^{2}$ (13)
$+\sup_{z\in\mathcal{Z}}\big\lVert\mathrm{IF}^{\sigma}(z,\sigma_{0},P)\big\rVert\cdot\frac{5\sqrt{2\ln(n)\ln{(2/\delta)}}}{\varepsilon n}\cdot U\bigg\}.$ (14)
@@ 3 Constructing CIs for the ATE under DP
Simply privatizing the variance to ensure DP is not sufficient to construct valid CIs in general: only in the asymptotic regime the intervals are valid, as no privatization is needed there and, as a result, the privatization noise equals zero. However, in finite samples, the privatized variance may underestimate the true variance, leading to undercoverage in the CIs.
Observe that the AIPW estimator $\hat{\tau}$ is a consistent estimator for the true ATE $\tau_{0}$ and asymptotically normally distributed with variance $\sigma^{2}_{\textrm{{AIPW}}}$ (e.g., Wager, 2024), i.e.,
@@
In finite settings, however, the privatization introduces additional noise on $\hat{\tau}$ (and thus variance), which we account for in the following when constructing our CIs. This is in line with prior literature postulating that DP inference should be conservative to ensure reliability even in small sample sizes (e.g., Karwa and Vadhan, 2017). Observe that $(\hat{\tau}_{\textrm{{DP}}}-\hat{\tau})=:B$ is itself a random variable (for fixed $\hat{\tau}$) following distribution $\mathcal{N}\left(0,r(\varepsilon,\delta,n)^{2}\right)$, where
@@
To incorporate the finite sample variance stemming from the privatization, we see from (17) that, for the variance $V$, it must hold that: $V=\sigma_{\textrm{{AIPW}}}^{2}+nr(\varepsilon,\delta,n)^{2}$.
@@
[…]
@@ 5 Experiments
Performance metrics: We evaluate PrivATE based on the notion of faithfulness (e.g., Schröder et al., 2024), i.e., we compute the empirical coverage of the CIs and assess whether it surpasses the $1-\alpha$ threshold.
@@ 5.1 Evaluation on synthetic datasets
Performance across varying privacy budgets: We vary the privacy budget $\varepsilon$ in Figure 4. We observe that PrivATE behaves as expected: for increasing privacy budgets, the intervals become tighter, eventually converging to the non-private estimate. We note that privacy inevitably induces a trade-off regarding the width of the interval, which we discuss in Supplement D.
Performance across varying sample sizes: We further analyze the performance of PrivATE across different sample sizes $n$ (see Fig. 4). The results are as expected and confirm our theoretical insights: For larger $n$, the CIs become tighter, converging to the asymptotic solution. Additionally, we observe that PrivATE is not sensitive to the choice of the base learner: both versions with (i) kernel ridge regression (=Kernel) and (ii) neural network (=NN) lead to similar results. This again demonstrates the flexibility of our framework.
@@ 5.2 Evaluation on medical datasets
We demonstrate the real-world applicability of PrivATE in a medical context on the Right Heart Catheterization (RHC) dataset (Connors et al., 1996). The dataset contains observations of 5735 patients admitted to intensive care units. We aim to predict the ATE of receiving a Swan-Ganz catheterization on the probability of death following hospitalization. Further details on the RHC dataset are in Supplement C.
@@
Results: In Table 2, we present the CIs returned by PrivATE compared to an adapted version of the naïve CIs constructed with a valid variance, which includes the correction term for the finite-sample variance stemming from ATE privatization. Again, we do not aim to benchmark the methods as the naïve CIs are non-private. Rather, we aim to understand whether the CIs from PrivATE are faithful. We observe that the PrivATE CIs include the valid, naïve CIs, which indicates faithfulness. In other words, our PrivATE gives CIs that remain informative while fulfilling $(\varepsilon,\delta)$-DP.
Medical interpretation: We observe that we always receive a positive private treatment effect $\hat{\tau}_{\textrm{{DP}}}$, indicating an average increase in the probability of death after receiving a Swan-Ganz catheter. The negative effect on survival is likely to be attributed to the life-threatening complications and risks of introducing the catheter, such as pulmonary artery rupture (e.g., Kearney and Shabot, 1995). The positive $\hat{\tau}_{\textrm{{DP}}}$ also coincides with the non-private ATE estimate $\hat{\tau}=1.17$ and is in line with findings in the literature (e.g., Connors et al., 1996; Sandham et al., 2003).
Extension to interventional RCT data: Our PrivATE framework applies to both observational and experimental data. In experimental randomized controlled trials (RCTs), the propensity score is known and thus does not need to be estimated. Note that in some cases, however, estimating the propensity score, even in RCTs, can be beneficial (Su et al., 2023). So far, we have evaluated PrivATE on observational data. To show the applicability of PrivATE to RCTs, we conduct additional experiments in Supplement E.
@@ A.1 Estimating the average treatment effect
Under the three standard assumptions of causal inference, i.e., positivity, consistency, and unconfoundedness (e.g., Curth and van der Schaar, 2021; Rubin, 2005), the ATE is point-identified as
@@
However, estimating $\hat{\tau}$ via the empirical average of the stated difference is only unbiased if the treatment assignment is randomized and if it does not depend on $X$ (i.e., if for a constant $\pi(x)$). To account for a non-random treatment assignment in observational data, the inverse propensity weighted (IPW) estimator constructs $\hat{\tau}$ by estimating $\hat{\pi}$ in the first step and then employing the estimates to take the propensity weighted empirical average of the outcomes (Rosenbaum and Rubin, 1983).
Yet, misspecification of $\hat{\pi}$ or $\hat{\mu}$ can lead to biased ATE estimates. As a remedy, the augmented inverse propensity weighted (AIPW) estimator (Robins et al., 1994) employs an orthogonal loss to fulfill the double robustness property. Doubly robust estimators are consistent if either $\hat{\mu}$ or $\hat{\pi}$ are correctly specified. The AIPW estimator is then given by
@@
Hence, in contrast to the IPW estimator, the AIPW estimator is an unbiased estimator of ATE and is asymptotically normal distributed (Wager, 2024). This is important as we (i) aim to consistently estimate the ATE, and (ii) can derive valid CIs due to the asymptotic normality to the estimator. In our work, we thus aim to derive differentially private CIs for the ATE estimated through the AIPW estimator.
@@ A.2 Differential privacy
In our work, we employ output perturbation based on the Gaussian mechanism. Of note, output perturbation is especially suitable for our task of constructing CIs for the estimated ATE due to our three reasons: output perturbation (i) retains the adherence to the causal assumptions, (ii) is model agnostic, and (iii) preserves the ability of the AIPW estimator to address the fundamental problem of causal inference (Schröder et al., 2025). In contrast, other approaches (i.e., input, objective, and gradient perturbations) fail to fulfill at least one of the points (i)–(iii) and would thus be not suitable for estimating the ATE with DP guarantees in a model-agnostic way.
@@
Lemma A.1 (Sequential composition property (Dwork et al., 2016)).
@@
Lemma A.2 (Post-processing property (Dwork, 2006)).
@@ A.3 Extended related work
Uncertainty quantification for causal quantities: Many different approaches for quantifying uncertainty in causal estimates exist in the literature. Many works rely on Bayesian methods (e.g., Alaa and van der Schaar, 2017; Hess et al., 2024; Jesson et al., 2020; Horii and Chikahara, 2024). However, Bayesian methods require prior distributions informed by domain knowledge. Therefore, they are less robust to model misspecification and, therefore, not suitable for our model-agnostic framework. A different stream focuses on estimating the conditional distribution of the treatment effect, which then can be employed for bootstrapping to construct uncertainty intervals (Ma et al., 2024; Melnychuk et al., 2023, 2024). Another stream of literature provides finite-sample uncertainty guarantees based on conformal prediction methods (e.g., Lei and Candès, 2021; Schröder et al., 2024). These methods can only construct intervals for the potential outcomes, not the treatment effects. Although one could construct from conformal prediction intervals for treatment effects based on the potential outcomes, those intervals commonly tend to be very wide and, therefore, less informative in practice. Importantly, none of the works above can provide $(\varepsilon,\delta)$-differentially private confidence intervals, which is the novelty of our work.
@@ A.4 Doubly robust variance estimation
Standard variance estimators for the AIPW estimator are, in general, only consistent if both nuisance functions are correctly specified (Gruber and van der Laan, 2012). Thus, these estimators are not doubly robust and, therefore, can easily lead to undercoverage of the CIs in practice, a highly undesirable scenario in safety-critical applications. We thus aim to estimate $\hat{\sigma}_{\textrm{{AIPW}}}^{2}$ through a doubly robust variance estimator.
A common such estimator is built upon nonparametric bootstrapping (Funk et al., 2011; Shook-Sa et al., 2024). However, bootstrapping requires many queries to the training data, which is highly disadvantageous for estimating the variance with DP guarantees. Further, bootstrapping is computationally expensive. This is unlike doubly robust variance estimators, which have been proposed only for specific applications, namely, survey data with missing data imputation (e.g., Haziza and Picard, 2012; Kim and Haziza, 2014).
In our work, we estimate $\hat{\sigma}_{\textrm{{AIPW}}}^{2}$ via the empirical sandwich estimator (Huber, 1967). This has the following benefits: (i) the empirical sandwich estimator is doubly robust, and (ii) it consistently estimates the variance of the AIPW estimator (Shook-Sa et al., 2024).
@@
[…]